### v1

# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

Abstract
A common dilemma while photographing a scene is whether to capture it in wider angle, allowing more of the scene to be covered but in lesser details or to click in narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a GAN-based model to learn to extract the visual quality parameters from a narrow angle shot and to transfer these to the corresponding wide-angle image of the scene. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
keywords:
Ultra-wide Image , Super Resolution , Generative Adversarial Network
[…]
[auth1]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
[…]
[auth2]organization=Department of Computer Science and Engineering, Jahangirnagar University, city=Dhaka
[…]
[auth3]organization=Department of Computer Science and Engineering, University of Dhaka, city=Dhaka
1 Introduction
Ultra-wide photography is a popular photographic trend in today’s world. This form of image-taking, though accommodating a greater field of view and thus offering a unique aesthetic, degrades the finer details better retained by the primary lens. This trade-off between the field of view covered and the quality of the finer details retained by the captured image places the users in an unwanted dilemma where one has to compromise the field of view for the details or vice-versa. Our primary goal in this paper is to enable the camera users to capture wide-angle images with quality that is identical to the narrow-angle shots while covering the original wide field of view.
The fundamental idea of capturing wide-angle images that retain the finer details associated with narrow-angle shots is to capture the scene of interest with two different lenses, both narrow and wide. The expectation here is that the narrow-angle shot provides the expected resolution while the wide-angle one carries the contents of the scene not covered in the narrow-angle counterpart. The idea would then be to render the contents captured by the wide-angle lens in the resolution that is represented by the narrow-angle counterpart. As it has been shown by the paper on Neural Style Transfer [6], the style and content of an image are separable and the stylistic representation of an image can be transferred into another image. We have leveraged this finding to transfer the resolution qualities of the narrow-angle shot into the wide-angle counterpart. Furthermore, we have implemented a pipeline that incrementally transfers the qualitative representations from the narrow-angle shot to the immediate wide-angle shot and repeats the process until the widest available image (often 0.5x or ultra-wide image) is reached. This involves capturing the scene of interest using all available lenses and incrementally applying the finest available quality (from the image with the narrowest field of view) through the stack of available images up to the image with the widest field of view.
In this paper, we propose a GAN-based architecture where the generator network produces realistic-looking enhanced versions of the wide FoV image while the discriminator component attempts to identify generated high-resolution results from the ground truth. In the first step, the spatial resolution of the original wide FoV image is upscaled using bicubic interpolation so that more details can be physically accommodated. Once upscaled, both the enhanced wide (from now onwards to be referred to as the wide FoV) and the narrow FoV images are partitioned into smaller patches so that visually similar patches may be utilized to assist the process of resolution transfer, inspired by the paper on Vision Transformer [11]. For the purpose of resolution transfer, we propose a modified version of the Gram matrix representation of an image, as used in the paper on Neural Style Transfer. Upon transferring the resolution from the narrow FoV patches into that from the wide FoV image, the synthetic image is passed through the discriminator setup that refines the seams between the tiles and the overall perceptual fidelity of the image, leveraging the task-specific loss functions we have designed. Finally, we perform this routine of resolution transfer throughout the stack of all available lenses to attain the highest possible resolution for the widest available field of view. Our contributions can thus be summarized into the following points,
  • 1.
    An end-to-end pipeline capable of transferring finer quality details to the wide-angle image from the corresponding narrow-angle image
  • 2.
    The novel concept of transferring visual quality across the stack of all available lenses
  • 3.
    An innovative modification of the Gram matrix that is able to capture the visual essence of an image
2 Related Works
Enhancing image quality has remained a central pursuit in computer vision. Earlier methods were primarily centered around super-resolution or upscaling techniques [1], which aimed to reconstruct high-resolution images from their low-resolution counterparts. While effective in producing sharper visuals, these models often introduced artifacts like oversmoothing, leading to a loss in texture and fine details. Later, generative adversarial models such as SRGAN [3], ESRGAN [2], and Real-ESRGAN [4] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs. However, these models still struggle with preserving structural consistency, especially in scenes where the narrow and wide FOV images exhibit differences in perspective or content.
Another stream of relevant research comes from neural style transfer. Initially developed for artistic applications [6], these models demonstrated that image attributes could be transferred from one image to another. Subsequent work improved efficiency using feed-forward networks [20] and adaptive normalization techniques [8], broadening their use cases. However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [24] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks.
Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [25] integrates shift convolutions with group-wise multi-scale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. ELAN’s architecture achieves a balance between performance and efficiency, making it suitable for applications requiring real-time processing. Further exploring attention-based models, DLGSANet [23] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multi-level Feature Fusion Network (DMFFN) [26] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. By employing two-level nested residual blocks with asymmetric structures, DMFFN efficiently fuses multi-level features, leading to improved edge preservation and faster convergence during training.
These contemporary models demonstrate significant progress in balancing performance and efficiency, particularly in scenarios requiring the enhancement of wide FOV images by leveraging information from narrow FOV counterparts. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Moreover, the integration of perceptual metrics such as LPIPS [16] has become increasingly prevalent, offering a more nuanced assessment of visual similarity that aligns closely with human perception. Incorporating human-in-the-loop evaluations further ensures that the enhanced images meet real-world expectations for quality and accuracy.
In this work, building upon these advancements, we aim to improve wide FOV image quality by transferring visual fidelity from narrow FOV versions of the same scene. By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
3 Method
In this section, we have described in detail the adopted methodologies and architecture of the model used to transfer high-quality details from the narrow-angle image into the wide-angle counterpart.
Figure 1: Model Architecture
In the case of training the model, we consider the original image as the wide FoV ground truth and extract the central portion to serve as the high-definition narrow FoV. The original wide FoV is then downsampled to simulate the captured wide FoV image.
3.1 Content Reconstruction
Our primary goal is to reconstruct the contents of the wide-angle image applying the visual qualities of the narrow-angle counterpart. By visual qualities we indicate the parameters that impact the perceptual cues which are color, brightness, contrast, and sharpness. For reconstructing the contents of the wide-angle image, we utilize the pixel-wise content loss. We predict the pixel entries for a white-noise image such that the following content loss is minimized.
$L^{(l)}_{content}(p,q)=\frac{1}{2}\sum_{i,j}(Q^{l}_{ij}-P^{l}_{ij})^{2}$
Here $p$ and $q$ are the original and white-noise images respectively and $P^{l}_{ij}$ and $Q^{l}_{ij}$ are the feature representations in the $l^{th}$ layer for $i^{th}$ filter at the position $j$.
3.2 Visual Quality Reconstruction
As mentioned earlier, the visual qualities are obtained from the narrow-angle image. The qualities are extracted from the image using the gram matrix approach. We have divided the image into a fixed number of patches and computed Gram matrices for each patch. The motivation to adopt this approach comes from the intuition that visually similar patches may exist between the narrow and the wide-angle images. Obtaining visual qualities from the patch of narrow-angle image that visually aligns with a patch of the wide-angle image would produce better results. Hence, we divide the entire narrow and wide-angle images into a predefined number of patches. For each patch of the wide-angle image, the patch from the narrow-angle image that is visually most similar is selected. The visual similarity score between patches is determined using a perceptual loss function described in the following section. Patches are taken to be visually similar if the similarity score exceeds a threshold. In case no patch from the narrow-angle image is found to exceed the threshold, a gram matrix representation of the entire narrow-angle image is used as a backup.
Figure 2 illustrates partitioning the image into fixed-sized patches. Images partitioned into fixed-sized patches just for illustrative purposes do not indicate the actual number of patches or their proportions used in the implementation. The patches are bound by black lines. Arrows and bounding boxes of different colors denote patches in the narrow FoV image that are visually similar to certain patches from the wide FoV image.
Figure 2: Image partitioning and patch similarity between narrow and wide FoV images.
3.2.1 Dividing into Patches
The narrow and wide-angle images are first divided into patches of dimension $128\times 128$. This particular size for the patches has been chosen to create a balance between structural information contained within an image and the computational cost incurred to process the patches. The idea of dividing the images into patches also takes inspiration from the paper on Vision Transformer [11].
3.2.2 Patch-wise Similarity
We use perceptual loss [17] to quantify the visual similarity between patches from narrow and wide-angle images. We use the mid layers of a VGG19 network to numerically determine the perceptual loss between the pair of patches. Following function is used to determine the loss,
$L_{feature}(p_{1},p_{2})=\frac{1}{C_{j}*H_{j}*W_{j}}||\phi_{j}(p_{2})-\phi_{j}(p_{1})||_{2}^{2}$
Here $p_{1}$ and $p_{2}$ are the patches from wide and narrow-angle images, $\phi_{j}(p_{1})$ and $\phi_{j}(p_{2})$ denote the feature map obtained from the $j^{th}$ layer, and $C_{j},H_{j}$ and $W_{j}$ are the number of channels, height, and width of the patches.
3.3 Generator Architecture
Figure 3: Generator Architecture
We utilize a ResNet [27] based generator with residual blocks. The input layer is followed by a collection of residual blocks to facilitate earlier layers in learning complex functions. The residual block relays the data into the upsampling block that generates the enhanced patch and passes through a final convolution layer to adjust the channel dimension.
3.3.1 Feature Map Construction
We consider the following parameters to effectively represent the visual qualities of an image (patch of an image in our case); $color,contrast,brightness,sharpness$. Considering each of these parameters as distinct features of a patch, an enhanced feature map representation of the patch is computed. The enhanced feature map is an augmentation of the individual features extracted from the patch and is formulated as follows,
$\hat{F_{l}}=[F_{l},F_{l_{color}},F_{l_{brightness}},F_{l_{contrast}},F_{l_{sharpness}}]$
Where $F_{l}\epsilon\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the feature map extracted using a VGG backbone network. Rest of the features are computed as follows,
  1. 1.
    Color: An average of the RGB channels forms an embedding that represents the overall intensities. Computed as,
    $F_{l_{color}}=\frac{1}{H_{l}W_{l}}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{j=1}F_{l}$
  2. 2.
    Brightness: Mean pixel intensities over all pixels are taken to indicate the perceived luminance of the patch. Computed as,
    $F_{l_{brightness}}=\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}F_{l}$
  3. 3.
    Contrast: Variance of the pixel intensities is computed to represent the variation of intensities, that is the contrast, within the patch. Calculated as,
    $F_{l_{contrast}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum^{C_{l}}_{i=1}\sum^{H_{l}}_{i=1}\sum^{W_{l}}_{k=1}(F_{l}-\mu_{l})^{2}}$
    here $\mu_{l}$ is the average intensity.
  4. 4.
    Sharpness: Sharpness is captured using the Laplacian filter. The Laplacian filter is multiplied to the feature map as follows,
    $F_{l_{sharpness}}=F_{l}*\bigtriangledown^{2}$
3.3.2 Gram Matrix Computation
Once the feature map is constructed, we can use this representation to find the gram matrix representation of the patch of the image for the current layer. The computation can be performed as follows,
$G_{l}=\hat{F}_{l}\hat{F}_{l}^{\top}$
The gram matrix is the product of the enhanced feature map to its transpose, essentially capturing the pair-wise correlation between features.
3.3.3 Loss Calculation
Visual quality loss is defined as the squared distance between the gram matrix representation of the source and target images. In our case, the source is the patch from the narrow FoV image while the target is a white-noise image which is eventually going to ”be” the resultant image by minimizing the loss defined above. The numerical representation of the loss is,
$L^{(l)}_{visual}=\sum_{l}w_{l}||G_{l_{q}}-G_{l_{p}}||^{2}$
Here $G_{l_{q}}$ and $G_{l_{p}}$ are the gram matrix representations of the patches from the white-noise and the narrow FoV images respectively. Mentionable that in case no particular match is found for a certain wide FoV patch in the narrow FoV, $G_{l_{p}}$ would denote the gram matrix representation of the entire narrow FoV image.
Hence the total generator loss is quantified as,
$L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$
3.4 Discriminator Loss
For the generated output image to be realistic and visually plausible, the results generated by the deconvolution network is passed through a discriminator component that aims to differentiate synthetic images from the real ones. We have designed two loss functions to assist the discriminator in this task. The losses are described as follows,
3.4.1 Seam Consistency Loss
This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
$L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
3.4.2 Perceptual Loss against Ground Truth
To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
Total discriminator loss is defined as follows,
$L^{(l)}_{D}=L^{(l)}_{seam}+L^{(l)}_{perceptual}$
3.5 Cascading Lens Stack
The above procedure is incrementally applied across the available lens array to render the widest possible image with a resolution comparable to the narrowest available shot of the scene. For example, if a camera has lenses with the following zoom capacities: $5x,3x,1x,0.5x$; visual qualities are iteratively transferred from the $5x$ to $3x$, then from $3x$ to $1x$ and so on until the widest available shot has been reached.
4 Experimentation
4.1 Datasets
We have used two datasets to ensure robust training of our proposed model.
  • 1.
    DIV8K [19]: High-resolution images widely used for super-resolution tasks. More than 1500 photos with a maximum resolution of 8K are included in the dataset. It covers a wide range of scene contents.
  • 2.
    Landscapes Dataset [18] (LHQ 1024×1024): With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes.
4.2 Generator
As stated earlier, the generator network takes in $128\times 128\times 3$ patches and performs an initial convolution followed by a ReLU layer. The feature map is then fed into an array of $64$ residual blocks. Residual blocks assist in preventing the gradient from dying out during back propagation. This is followed by $2$ upsampling blocks, responsible for physically upscaling the patch using Pixel Shuffle layers. The upsampled patches passed through the output convolution block to adjust the channel dimension and come out as enhanced patches with a dimension of $256\times 256\times 3$.
4.3 Discriminator
We use a ResNet18 based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scaler denoting the realness of the input image.
4.4 Simulating Narrow and Wide FoVs
A high-quality narrow FoV and two wide FoV (one low-quality and one high-quality) images are required for training. High-quality wide images are used to calculate the loss, whereas low-quality wide and high-quality narrow images are used as model inputs. The original images from the dataset are treated as high-quality wide FoV images as well as the ground truth. To simulate high-quality narrow and low-quality wide-angle images from single-image datasets we have used the following techniques.
High-quality narrow FoV simulation: For each image in the dataset, we extract a high-resolution central region with a zoom factor of $\frac{5}{3}$. This region is then embedded into a larger blank canvas, preserving the original central region’s quality. The resulting image replicates the characteristics of a narrow FoV while maintaining the high quality of the original.
Low-quality wide FoV simulation: For the DIV8K dataset with 8K resolution, we simulated low-quality wide FoV images by first downsampling each image to 4096 × 4096 and then upsampling it back to 8K. Similarly, for the Landscapes dataset (LHQ 1024 × 1024), we downsampled each image to 512 × 512 and then upsampled it back to 1024 × 1024. This intentional degradation process mimics low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions. The description of workflow provided in this paper is in light of the Landscape dataset, that is, assuming the original image to be 1024 x 1024. The image reconstruction pipeline has been shown in the following Figure-4.
Figure 4: Enhanced image reconstruction pipeline
4.5 Training Details
Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss. The full generator-discriminator framework was then trained for an additional 200 epochs, totaling 400K iterations. The generator consists of 16 residual blocks, each with two 3×3 convolutional layers, batch normalization, and parametric ReLU activation, allowing effective feature extraction while keeping training stable. To enhance high-frequency details, two upsampling layers using pixel shuffle progressively increase the resolution. The final 3×3 convolutional layer reconstructs the high-resolution image.
The convergence behavior of the visual quality loss for a single patch in the generator network is shown in Figure 5. It highlights the effectiveness of our proposed enhanced Gram matrix, which causes an earlier convergence by approximately 30K iterations compared to the traditional approach. This earlier stabilization suggests that our method improves the efficiency of optimization, reducing the training time required while maintaining reconstruction quality.
Figure 5: Faster convergence of visual quality loss with the enhanced Gram matrix.
To further stabilize the training and prevent mode collapse, spectral normalization was applied to the discriminator while gradient clipping controlled exploding gradients. The model was optimized using a combination of content loss (L1), perceptual loss (VGG-based), and adversarial loss, ensuring a strong balance between detail accuracy and perceptual quality.
4.6 Experiment Results
Table 1 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
Table 2 illustrates the cascading effect of the outputs derived from images captured at different zoom levels. In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output. The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image. This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
Narrow FoV Input wide FoV Our result
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Experimental results showing the output generated by our model when processing narrow and wide input images.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2x Zoom 1x Zoom 1x Output
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
1x Output 0.5x Zoom 0.5x Output
Table 2: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS[16] metrics over Set5[12], Set14[13] and BSD100[14] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN[2], LatticeNet[24], ELAN[25], DLGSANet[23] and DMFFN[26] in Table 3 and Table 4.
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 32.73 32.18 32.43 32.46 32.41 32.80
Set14 28.99 28.61 28.78 28.79 28.74 30.37
BSD100 25.16 27.57 27.69 27.70 27.74 26.58
Table 3: PSNR Comparison (highest marked in red)
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 0.9011 0.8943 0.8975 0.8993 0.8981 0.9118
Set14 0.7917 0.7812 0.7858 0.7871 0.7874 0.7806
BSD100 0.7455 0.7355 0.7406 0.7415 0.7421 0.8261
Table 4: SSIM Comparison (highest marked in red)
We have presented our comparison with bicubic interpolation and Multiscale GAN [15] on the LPIPS metric in Table 5. It is to be noted that LPIPS values are within the range $[0,1]$, with a lower value denoting the generated image is perceptually more aligned with the ground truth.
Dataset bicubic Multiscale GAN Ours
Set5 0.5657 0.2823 0.0795
Set14 0.6114 0.3717 0.2032
BSD100 0.5726 0.4112 0.3361
Table 5: LPIPS Comparison (best marked in red)
Figure 6 presents a comparative analysis of our method against several prior super-resolution models, including bicubic interpolation, LatticeNet [24], ELAN [25], and DLGSANet [23]. The leftmost image represents the original ground truth, with the red bounding box highlighting the region of interest for focused evaluation. The comparative results illustrate variations in texture reconstruction, detail preservation, and structural fidelity across different approaches. Notably, our method, displayed in the bottom-right corner, achieves superior sharpness and consistency, effectively reducing blurring and artifacts observed in other models.
Figure 6: Visual comparison of super-resolution models, including bicubic interpolation, LatticeNet, ELAN, and DLGSANet. The red bounding box highlights the region of interest, with our method (bottom-right) achieving superior detail preservation and structural fidelity
6 Conclusion
We proposed a GAN-based approach to enhance wide-angle images by transferring fine details from a narrow-angle shot. Our method balances coverage and detail through multi-zoom image fusion, leveraging perceptual losses and an improved Gram matrix representation. The cascading lens stack further extends its application to multi-lens imaging systems.
Extensive experiments showed that our approach outperforms existing super-resolution and image synthesis methods, achieving superior PSNR, SSIM, and LPIPS scores while reducing artifacts. This opens new possibilities for computational photography, smartphone imaging, and digital zoom. Future work will focus on improving performance under extreme lighting, dynamic scenes, and real-time constraints, with potential enhancements through self-supervised learning.
During the preparation of this work, the author(s) used ChatGPT and Quillbot in order to ensure grammatical correctness of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.