### v2

# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

Abstract
A common dilemma while photographing a scene is whether to capture it at a wider angle, allowing more of the scene to be covered but in less detail or to click in a narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a Generative Adversarial Network (GAN)-based model to learn to extract the visual quality parameters from a narrow-angle shot and to transfer these to the corresponding wide-angle image of the scene using residual connections and an attention-based fusion module. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
keywords:
Ultra-wide Image, Super Resolution, Generative Adversarial Network, Attention-based Fusion
1 Introduction
Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame. While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses, as discussed in the paper [2]. Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view. In this work, we seek to resolve this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
Our approach hinges on jointly leveraging both narrow- and wide-angle captures of the same scene. The narrow-angle image supplies high-resolution texture and detail, while the wide-angle image carries context beyond the primary lens’s range. We introduce an end-to-end pipeline that incrementally transfers visual fidelity from the finest (narrowest) view through to the widest view in a multi-lens stack. First, we partition both the wide image and the corresponding narrow-angle image into spatially aligned patches inspired by the paper on Vision Transformer [3]. Each patch pair is fed into a generator, which fuses content and detail via cross-view attention and further super-resolves the wide patch using pixel-shuffle layers—progressively enhancing resolution and texture across the full frame.
Central to our method is a novel GAN [4] architecture: a generator tailored to fuse cross-view content and detail features through attention mechanisms and a discriminator designed to enforce both global realism and seamless patch blending. We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity. Applying this procedure iteratively across all available zoom levels—from telephoto to ultra-wide—we end up with a single image that combines the sweeping perspective of the widest lens with the crisp detail of the primary lens. Our key contributions are
  • An end-to-end pipeline that transfers high-resolution detail from narrow to wide-angle images.
  • A GAN framework with cross-view attention and an enhanced Gram matrix formulation for robust detail propagation.
  • A cascading lens-stack strategy that scales seamlessly from telephoto to ultra-wide captures.
2 Related Works
Enhancing image quality has remained a central pursuit in computer vision. Earlier methods were primarily centered around super-resolution or upscaling techniques [5], which aimed to reconstruct high-resolution images from their low-resolution counterparts. While effective in producing sharper visuals, these models often introduced artifacts like oversmoothing, leading to a loss in texture and fine details. Later, generative adversarial models such as SRGAN [6], ESRGAN [7], and Real-ESRGAN [8] improved perceptual realism by emphasizing texture fidelity and producing visually convincing outputs. However, these models still struggle with preserving structural consistency.
Another stream of relevant research comes from neural style transfer. Initially developed for artistic applications [9], these models demonstrated that image attributes could be transferred from one image to another. Subsequent work improved efficiency using feed-forward networks [10] and adaptive normalization techniques [11], broadening their use cases. However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
Figure 1: Overall model architecture
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [12] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks. Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [13] integrates shift convolutions with group-wise multiscale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. The ELAN architecture achieves a balance between performance and efficiency, making it suitable for applications that require real-time processing.
Further exploring attention-based models, DLGSANet [14] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multilevel Feature Fusion Network (DMFFN) [15] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. Using nested two-level residual blocks with asymmetric structures, DMFFN efficiently fuses multilevel features, leading to improved edge preservation and faster convergence during training. These contemporary models demonstrate significant progress in balancing performance and efficiency. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Perceptual metrics such as LPIPS [16] have become widely adopted for evaluating visual similarity in a way that aligns closely with human perception. Beyond perceptual loss, incorporating visual cues that guide perception—namely color, brightness, contrast, and sharpness—is equally important. We select these attributes based on extensive literature: [17] identifies them as key appearance properties in image quality assessment; [18] highlights the role of color in quality estimation; [19] emphasizes the critical balance of brightness and contrast in enhancement; and [20] combines contrast and sharpness to assess enhancement reliability.
By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
3 Method
The core philosophy of our solution is to use a cross-view attention mechanism to fuse content and visual quality features from a given pair of wide and narrow-angle images. The images are initially partitioned into fixed-size patches. We pick a pair of wide and narrow patches from the pool and pass them across the network architecture, as described below.
3.1 Image Partitioning
Both narrow and wide FoV images are partitioned into $128\times 128$-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data. Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [21] and stored in the pool, ready to be pushed into the generator network. A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [22] backbone is used to determine the perceptual distance between patches.
Figure 2: Image partitioning and patch similarity between narrow and wide FoV images.
Figure 3: Detailed architecture for the Generator component
3.2 Network Architecture
The architecture of the generator network is built upon a stack of ResNet-like [23] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool. The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches respectively. Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps respectively. The pipeline consists of the following major components,
3.2.1 Structural Feature Encoder
Wide FoV patches are encoded with traditional CNN-based encoders to capture the content details. Encoded wide patch represented as $E_{w}$
3.2.2 Visual Feature Encoder
These encoders are specialized in extracting the visual parameters from a patch and are used to encode the narrow FoV patches. We use the Gram matrix approach to represent the visual features of a patch and include the following additional parameters in the Gram matrix representation: color, brightness, contrast, and sharpness. The feature map is constructed as follows,
  • The original feature map $F_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is generated using a VGG-19 network.
  • Color information for each channel over spatial dimensions, also referred to as global average pooling [24], is computed as:
    $[F_{l_{\text{color}}}]_{c}$ $=\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w),$
    for $\quad\text{for }c=1,\dots,C_{l}$
  • Global mean pixel intensity across all channels and spatial dimensions indicates the perceived luminance (brightness) of the patch. Computed as:
    $F_{l_{\text{brightness}}}=\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w)$
  • We represent the overall contrast of the patch using the standard deviation:
    $F_{l_{\text{contrast}}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}\left(F_{l}(c,h,w)-\mu_{l}\right)^{2}}$
    where $\mu_{l}=F_{l_{\text{brightness}}}$ is the global mean defined above.
  • Sharpness is accounted for using the Laplacian operated, as mentioned in this paper [25]:
    $F_{l_{\text{sharpness}}}(c,h,w)=(F_{l}(c,\cdot,\cdot)*\nabla^{2})(h,w)$
    where * denotes the convolution and $\nabla^{2}$ is the Laplacian kernel.
  • Finally, we augment the original feature map obtained from VGG-19 as follows:
    $\hat{F}_{l}=[F_{l},F_{l_{\text{color}}},F_{l_{\text{brightness}}},F_{l_{\text{contrast}}},F_{l_{\text{sharpness}}}]$
    where scalar or vector features are broadcast to match the spatial dimensions of $F_{l}$.
Once the enhanced feature map has been computed, we perform a dot product between the enhanced feature map and its transpose to attain the Gram matrix representation $E_{n}$, carrying the visual essence of the narrow FoV patch.
$E_{n}=\hat{F}_{l}\hat{F}_{l}^{\top}$
3.2.3 Cross-Attention Residual Blocks
Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch.
  • Encoded wide FoV patch $E_{w}$, representing the content data, acts as the query matrix
  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices
  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed using the scaled dot product equation introduced in the attention paper [26],
    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{k}}})E_{n}$
  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition, producing the feature map representation of the enhanced wide FoV patch, $E$
    $E=E_{w}+Attention(E_{n},E_{w})$
3.2.4 Upsampling Layers
The output of the final residual block is passed through a pixel shuffle [27] layer with a scale factor equal to 2 to super-resolve the patch, denoted by $E^{*}$.
3.3 Loss Calculation
Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth. Then the patches are assembled using seamless cloning, to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator which computes a seam consistency loss over the entire image. The losses are described as follows,
3.3.1 Generator Loss
  1. 1.
    Content Loss: We find the deviations of the contents of the generated patch and ground truth with the following pixel-wise difference,
    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{i}(E^{o}_{w}-E^{g}_{w})^{2}$
    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches.
  2. 2.
    Visual Quality Loss: This loss is calculated as the difference of the Gram matrix representation of the generated patch and the narrow FoV patch using the following equation,
    $L^{(l)}_{visual}=\sum_{l}w_{l}||E^{o}_{n}-E^{g}_{n}||^{2}$
    Here $E^{g}_{n}$ and $E^{o}_{n}$ are the gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively.
  3. 3.
    Total generator loss can be written as,
    $L^{(l)}_{G}=L^{(l)}_{content}+L^{(l)}_{visual}$
3.3.2 Discriminator Loss
For the generated output image to be realistic and visually plausible, the results generated by the deconvolution network is passed through a discriminator component that aims to differentiate synthetic images from the real ones. We have designed two loss functions to assist the discriminator in this task. The losses are described as follows,
  1. 1.
    Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
    $L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
  2. 2.
    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
    Total discriminator loss is defined as follows,
    $L^{(l)}_{D}=L^{(l)}_{seam}+L^{(l)}_{perceptual}$
3.4 Cascading Lens Stack
The above procedure is incrementally applied across the available lens array to render the widest possible image with a resolution comparable to the narrowest available shot of the scene. For example, if a camera has lenses with the following zoom capacities: $5x,3x,1x,0.5x$; visual qualities are iteratively transferred from the $5x$ to $3x$, then from $3x$ to $1x$ and so on until the widest available shot has been reached.
4 Experimentation
4.1 Datasets
We have used Landscapes Dataset [28] (LHQ 1024×1024) to train the proposed model. With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes. We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed into the training loop in the ascending order of pixel variance [29], to ensure “easier” images are encountered first.
Figure 4: Convergence of loss curves shown over multiple iterations
[H] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
Table 1: Experimental results demonstrating the cascading effect of multi-zoom image upscaling.
4.2 Simulating Narrow and Wide FoVs
High-quality narrow FoV simulation: For each image in the dataset, we extract a high-resolution central region with a zoom factor of $\frac{5}{3}$. This region is then embedded into a larger blank canvas, preserving the original central region’s quality. The resulting image replicates the characteristics of a narrow FoV while maintaining the high quality of the original.
Low-quality wide FoV simulation: We downsampled each image by applying a blur kernel and adding noise following this paper [30], as well as resizing to 512 × 512, and then upsampled it (using bicubic interpolation) back to 1024 × 1024. This intentional degradation process mimics the wide image of low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions.
4.3 Generator
The generator takes in a pair of patches, each of dimension $128\times 128\times 3$. The wide FoV patch is passed through a CNN encoder with a final batchNorm layer to produce the content feature map. The narrow FoV patch initially undergoes feature extraction through the $Conv2_{2},Conv3_{2},Conv4_{2}$ layers of a VGG-19 backbone. The feature map produced by VGG-19 is enhanced by the visual properties computed from the original patch. We use 10 attention-based residual blocks, each with two $3\times 3$ convolution units, a ReLU, followed by the cross-view attention score calculation unit, outputs of which are added to the wide FoV patch. The details-infused wide FoV patch is then passed through a series of convolution layers interleaved by pixel shuffle layers to physically upscale the patch to a dimension of $256\times 256\times 3$.
4.4 Discriminator
We use a ResNet18 [31] based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending [32]. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scalar denoting the realness of the input image.
4.5 Enhanced image reconstruction pipeline
Figure 5 highlights the steps undertaken to reconstruct the enhanced.
Figure 5: Enhanced image reconstruction pipeline
4.6 Training Details
Key hyperparameters for training and evaluation included a patch size of 128 × 128, Adam optimizer with a learning rate of 0.0001 with exponential decay every 10 epochs, a batch size of 16, perceptual loss layers (Conv2_2, Conv3_2, Conv4_2 of VGG19), and a threshold for patch similarity empirically set to 0.7.
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images. The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations.
The convergence behavior of the generator loss has been illustrated in Figure 4. These curves represent the first 15 epochs where the generator has been trained solo. Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [33] method.
Narrow FoV Input wide FoV Our result
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Experimental results showing the output generated by our model when processing narrow and wide input images.
[H] Dataset bicubic Multiscale GAN Ours Set5 0.5657 0.2823 0.0795 Set14 0.6114 0.3717 0.2032 BSD100 0.5726 0.4112 0.3361
Table 3: LPIPS Comparison (best marked in red)
Figure 6: Visual comparison of our results against the mentioned prior arts
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 32.73 32.18 32.43 32.46 32.41 32.80
Set14 28.99 28.61 28.78 28.79 28.74 30.37
BSD100 25.16 27.57 27.69 27.70 27.74 26.58
Table 4: PSNR Comparison (highest marked in red)
Dataset ESRGAN LatticeNet ELAN DLGSANet DMFFN Ours
Set5 0.9011 0.8943 0.8975 0.8993 0.8981 0.9118
Set14 0.7917 0.7812 0.7858 0.7871 0.7874 0.7806
BSD100 0.7455 0.7355 0.7406 0.7415 0.7421 0.8261
Table 5: SSIM Comparison (highest marked in red)
4.7 Experiment Results
Table 2 presents the results of our experiment, where a narrow and a low-resolution wide image are used as input. These images are processed through our model to generate the final output.
Table 1 illustrates the cascading effect of the outputs derived from images captured at different zoom levels. In the first row, the input images consist of a 2x (telephoto) and a 1x (standard) zoom, which collectively produce a high-resolution 1x (standard) output. The second row takes the output from the previous row as one input and combines it with a low-resolution 0.5x (wide) image, ultimately generating a high-resolution 0.5x (wide) image. This sequence effectively simulates how our method can be applied in a cascading manner using all available zoom lenses.
5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS [16] metrics over Set5 [34], Set14 [35] and BSD100 [36] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] in Table 4 and Table 5. We have presented our comparison with bicubic interpolation and Multiscale GAN [37] on the LPIPS metric in Table 3. Visual comparison of our method against prior super-resolution models has been presented in Figure 6.
6 Conclusion
In this work, we presented a novel approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene. The core of our method is a patch-based cross-view attention mechanism that selectively transfers fine-grained details from narrow-FoV patches to their wide-FoV counterparts, guided by Gram matrix-based visual encoding that captures texture statistics for accurate alignment. To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
We trained our model on the Landscape HQ dataset to ensure coverage of both structural diversity and fine detail. For evaluation, we used standard benchmark datasets including BSD100, Set5, and Set14. Experimental results demonstrate that our method consistently outperforms strong baselines such as ESRGAN, ELAN, DLGSANet, DMFFN, and LatticeNet in both PSNR and SSIM metrics. Qualitative comparisons further confirm the perceptual superiority of our results.
Overall, our framework successfully addresses the challenge of enhancing ultra-wide FoV images by combining cross-view attention, texture-aware encoding, and a cascading lens-based reconstruction strategy. Future extensions could explore its application to video enhancement and real-time inference scenarios.